EDITORIAL

Terry Goodman

I wanted to bring to your attention a very interesting article, “It’s Time To Tell the Kids: If You Don’t Do Well in High School, You Won’t Do Well in College (or on the Job).” The article, found in the Spring, 2004 issue of The American Educator, focuses on the role that students’ high school educations have in preparing them for college. The author, James E. Rosenbaum, begins by pointing out that as of 1992, 84 percent of high school seniors planned to get a college degree, but data from the high school classes of 1972, 1982, and 1992 showed that only 45% to 49% of students who enter college and earn more than 10 credits actually earn a bachelor’s degree. Further, for students with high school averages of C or lower, less than 50% of them will earn even one college credit.

Rosenbaum goes on to discuss three social transformations that have worked together to increase the percentage of students who want to go to college. These three transformations: the earnings advantage of college graduates has grown, college has become much more accessible, and virtually all community colleges and many 4–year institutions have adopted a policy of open admissions have contributed to six misconceptions.

1. College success is not linked to high school preparation.

2. College plans lead to increased school effort.

3. High school homework doesn’t matter for college success.

4. Going to college means taking college–level classes.

5. Going to college for a two– or four–year degree takes two or four years.

6. School counselors should not offer discouraging words about the hard work necessary for college success.

Data are presented that both document the presence of these misconceptions as well as refute their validity.

The article goes on to discuss how the “rules of the game” have changed. Among these new rules:

1. All students can plan to get a college degree; but if they are unprepared, they must be willing to repeat high school courses in college, spending extra time, money, and effort in non–credit, remedial courses.

2. College completion, as opposed to enrollment, requires increased high school effort.

3. Students whose college prospects are dim should be provided good information about alternatives to college.

Rosenbaum also emphasizes that the data show that if you take hard classes, do all your homework, and get good grades in high school, you will generally be academically ready for college. Of particular interest is the fact that the further a student goes in high school mathematics, the better his/her chances of earning a college degree. Of high school students who take calculus, 79.8% of them get a bachelor’s degree while only 23.1% of those who complete up through geometry and 39.5% of those who complete up through Algebra 2 get a bachelor’s degree.

Finally, Rosenbaum cites a study conducted by Adelman (1999) in which the differences between those high school students who did and did not graduate from college. Among these findings are the following:

1. Of all the indicators of academic preparation, the one that is the strongest is taking rigorous and intense courses in high school.

2. Taking rigorous and intense high school courses has a greater impact on African–American and Latino students than on white students.

3. Of all the high school courses, the highest level of mathematics taken is the most important for college success.

4. Academically well–prepared students are likely to graduate from college regardless of their social background. Unprepared students of all backgrounds are not likely to do so. I would encourage all our readers to read the Rosenbaum article. You can find it

online at: \underline{http : //www.aft.org/american educator/spring2004/tellthekids.html}.

Adelman, C. (2004). Principal Indicators of Student Academic Histories in Postsecondary Education, 1972–2000, Washington, D.C., U.S. Department of Education.